# Linear colorings of graphs

### v1

@@ 3 General bounds
Following [KOPS21], Bose et al. [BDH+22] relaxed the notion of a $k\times k$ grid to a family of graphs called $k\times k$ pseudogrids, that have the following properties:
@@
    A graph $G$ contains $\boxplus_{k}$ as a minor if and only if $G$ contains a $k\times k$ pseudogrid as a subgraph (that is, $k\times k$ pseudogrids are exactly the subgraph-minimal graphs containing $\boxplus_{k}$ as a minor).
@@
[[CNP21]] For every two integers $w,d>0$ and every graph $G$, if $\operatorname{\chi_{\mathrm{cen}}}(G)\geqslant\alpha wd$ then either $\operatorname{\mathbf{tw}}(G)\geqslant w$ or $G$ contains as a subgraph a subcubic tree $T$ with $\operatorname{\chi_{\mathrm{cen}}}(T)\geqslant d$.
@@
In the remaining outcome of Section 3, $G$ contains a subcubic tree $T$ with $\operatorname{\chi_{\mathrm{cen}}}(T)\geqslant d$ as subgraph. We apply Section 2.2 and get $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant\operatorname{\chi_{\mathrm{lin}}}(T)\geqslant d/\log 3=\left\lfloor\sqrt{k/\alpha}\right\rfloor/\log 3$. So similarly as above $k\leqslant c_{2}\operatorname{\chi_{\mathrm{lin}}}(G)^{2}$ for some constant $c_{2}>0$. Overall the claimed statement holds for $c=\max(c_{0},c_{1},c_{2})$. ∎
@@
$\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant c\cdot(\operatorname{\mathbf{tw}}(G)+1)\cdot\left(\operatorname{\chi_{\mathrm{lin}}}(G)+\frac{1}{\log 3}\right)$
@@
Treewidth does not increase when taking subgraphs, so only the second outcome of the theorem may hold. Hence $G$ contains a subcubic tree $T$ with $\operatorname{\chi_{\mathrm{cen}}}(T)\geqslant d$.
@@
As a corollary (and using that when $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 2$, $\operatorname{\chi_{\mathrm{cen}}}(G)=\operatorname{\chi_{\mathrm{lin}}}(G)$, see Section 5.1) we get the following bound.
@@
$\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant c\cdot(\operatorname{\mathbf{tw}}(G)+1)\cdot\operatorname{\chi_{\mathrm{lin}}}(G)$
@@
In graph classes where the treewidth is linearly bounded from above by the clique number (which is a lower bound for $\operatorname{\chi_{\mathrm{lin}}}$), Section 3 implies a quadratic bound. In particular, this is the case for intersection graphs of connected subgraphs of graphs of bounded treewidth. For a class $\mathcal{G}$ of graphs, let $I(\mathcal{G})$ be the class of region intersection graphs of graphs in $\mathcal{G}$, defined as follows. For $H\in\mathcal{G}$ and a family $\{H_{j}\}_{j\in J}$ of connected subgraphs of $H$, let $G$ be the graph with vertex set $J$ in which two distinct vertices $i$ and $j$ are adjacent if and only if $H_{i}$ and $H_{j}$ have a vertex in common. Region intersection graphs have been studied as a common generalization of many classes of geometric intersection graphs (see [Lee17]). For any graph class $\mathcal{G}$ of graphs with bounded treewidth, region intersection graphs of graphs in $\mathcal{G}$ have treewidth linearly bounded from above by the clique number. This follows from the following property, shown in the proof of [BGT98, Lemma 2.4],
@@
[[BGT98]] Let $k\geqslant 0$ be an integer and let $\mathcal{G}$ be a class of graphs with treewidth at most $k$. Then each graph $G\in I(\mathcal{G})$ has a tree decomposition in which each bag is a union of at most $k+1$ cliques.
@@
Let $k\geqslant 0$ be an integer and let $\mathcal{G}$ be a class of graphs with treewidth at most $k$. Then each graph $G\in I(\mathcal{G})$ satisfies $\operatorname{\chi_{\mathrm{cen}}}(G)=\operatorname{\mathcal{O}}(\operatorname{\chi_{\mathrm{lin}}}(G)^{2})$.
@@
Let $G\in I(\mathcal{G})$. By Section 3, $G$ has a tree decomposition in which each bag is a union of at most $k+1$ cliques. Hence, $\operatorname{\mathbf{tw}}(G)\leqslant(k+1)\omega(G)-1$. By Section 3, we infer that
$\operatorname{\chi_{\mathrm{cen}}}(G)=\operatorname{\mathcal{O}}((\operatorname{\mathbf{tw}}(G)+1)\cdot\operatorname{\chi_{\mathrm{lin}}}(G))=\operatorname{\mathcal{O}}((k+1)\omega(G)\cdot\operatorname{\chi_{\mathrm{lin}}}(G))=\operatorname{\mathcal{O}}(\operatorname{\chi_{\mathrm{lin}}}(G)^{2})\,.$
since $\omega(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)$ and the constant $k+1$ is subsumed by the $\mathcal{O}(\cdot)$-notation. ∎
Since every chordal graph is an intersection graph of subtrees in a tree (see [Bun74, Wal78, Gav74]) and every circular-arc graph is an intersection graph of paths in a cycle, the cases $k\in\{1,2\}$ of Section 3 imply the following generalizations of the analogous result from [KOPS21] for interval graphs (Section 1).
@@
For every tree $T$, $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant c\cdot\operatorname{\chi_{\mathrm{lin}}}(T)$ where $c=\alpha\log 3<3.7\,.$
@@
4 When centered is almost linear
Interesting cases of graphs satisfying Section 1 are the graphs for which the two chromatic numbers are equal. A sufficient condition for this is that every linear coloring of the graph is centered. Three graph classes with this property were already identified by Kun et al. [KOPS21]. In Section 4.1, we identify two more.
@@ 4.1 Graphs for which every linear coloring is centered
Kun et al. [KOPS21] observed that this property holds for graphs with independence number at most two, cographs, and graphs with maximum degree at most $2$. The result for graphs with independence number at most two can be generalized as follows.
@@
$\overline{H}$ is connected. Since $\overline{H}$ is paw-free, by a result of Olariu [Ola88] $\overline{H}$ is either triangle-free or complete multipartite. In the first case the independence number of $H$ is at most $2$, so by [KOPS21, Lemma 1] $\varphi$ is a centered coloring, a contradiction. In the second case, as $H$ is connected, $H$ is a complete graph and the same conclusion follows.
@@
$\overline{H}$ is disconnected. Then $H$ has a bipartition $(A,B)$ of its vertex set into two nonempty sets with all edges between them. As $\varphi$ is a proper coloring, $\varphi$ colors $A$ and $B$ with disjoint sets of colors. Besides, by Section 4.2, in one of $A$ and $B$ $\varphi$ assigns different colors to all the vertices. We may assume without loss of generality that the vertices of $A$ gets colored with different colors. Then, any vertex of $A$ is a center of $H$, contradicting the choice of $H$. ∎
A useful observation, also used in [KOPS21], is that the property holds for graphs such that every connected induced subgraph has a Hamiltonian path. {observation} If $G$ is such that every connected induced subgraph of $G$ has a Hamiltonian path, then every linear coloring of $G$ is centered. Indeed, if $H$ admits a linear coloring $\psi$, then for every connected subgraph $H$ of $G$, $V(H)$ induces a connected subgraph that has a Hamiltonian path $P_{H}$. Then $P_{H}$ admits a center thus $H$ admits a center, and therefore $\psi$ is a centered coloring of $G$.
A graph $G$ is an interval graph if it is an intersection graph of a family of closed intervals on the real line. If moreover none of the intervals contains another, then $G$ is a proper interval graph. Kun et al. [KOPS21] showed that every interval graph $G$ satisfies $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)^{2}$. If $G$ is a proper interval graph, then $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$, as implied by the fact that every proper interval graph is (claw, net)-free [^3] The claw is the graph obtained by attaching 3 pendant vertices to a vertex, and the net is the graph obtained by attaching a pendant vertex to each vertex of a $K_{3}$. and the following. {proposition}If $G$ is (claw, net)-free, then every linear coloring of $G$ is centered. In particular, $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$.
@@ 4.2 Complete multipartite graphs and complements of rook’s graphs
Given a positive integer $k$, a complete $k$-partite graph is any graph $G$ such that $V(G)$ admits a partition into $k$ nonempty parts such that two vertices of $G$ are adjacent if and only if they belong to different parts. A complete multipartite graph is any graph $G$ that is complete $k$-partite for some $k$. Note that if $G$ is complete multipartite, then $G$ is complete $k$-partite for a unique $k$; moreover, the corresponding partition into $k$ parts is unique. Complete multipartite graphs are cographs, hence, every linear coloring is centered (see [KOPS21]). In the next theorem we determine the exact values of the corresponding chromatic numbers. We start with a simple lemma that we use twice, so we extracted the statement.
{lemma}
@@
{theorem}
Let $G$ be a complete multipartite graph. Then $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$, with both quantities equal to $n-p+1$ where $n$ is the order of $G$ and $p$ is the maximum cardinality of a part of $G$.
Proof.
To see that $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant n-p+1$, Observe that we can color the largest part $Y$ of $G$ with the same color and all the other vertices with different colors. This is a centered coloring since every subgraph of $G$ either contains a single vertex from $Y$, or contains a uniquely colored vertex from $G-Y$.
To show that $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant n-p+1$, consider a linear coloring $\varphi$ and let $k$ be the number of colors used. If all vertices of $G$ are colored with unique colors, then $k=n\geqslant n-p+1$. Hence, we may assume that two vertices $u,v$ receive the same color. Then $u$ and $v$ are non-adjacent and thus belong to the same part $X$ of $G$. Since the pair $(X,V(G)\setminus X)$ is a bipartition of $V(G)$ with all edges between $X$ and $V(G)\setminus X$, by Section 4.2, $\varphi$ assigns different colors to all the vertices in one of $X$ and $V(G)\setminus X$. Since $\varphi(u)=\varphi(v)$, we infer that $\varphi$ assigns different colors to all the vertices in $V(G)\setminus X$. Furthermore, since $\varphi$ is a proper coloring, no color of a vertex in $V(G)\setminus X$ can be used on any vertex in $X$. Hence, $k\geqslant n-|X|+1\geqslant n-p+1$, as claimed. We conclude that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant n-p+1\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)$, therefore, equalities must hold throughout. ∎
For all $n,m\geqslant 1$, the $n\times m$ rook’s graph is a graph on $nm$ vertices, arranged into $m$ columns with $n$ vertices each and into $n$ rows with $m$ vertices each, such that two vertices are adjacent if and only if there are in the same column or in the same row.
@@
Assume now that $m\geqslant 2$ and $n\geqslant 3$, with $n\geqslant m$. Since $G_{n,m}$ is a subgraph of the complete multipartite graph with $m$ parts of $n$ vertices, we obtain from Sections 2.2 and 4.2 that $\operatorname{\chi_{\mathrm{cen}}}(G_{n,m})\leqslant nm-n+1$; so let us show that $\operatorname{\chi_{\mathrm{lin}}}(G_{n,m})\geqslant nm-n+1$. Suppose for a contradiction that there exists a linear coloring $\varphi$ of $G_{n,m}$ with at most $nm-n$ colors.
@@
Suppose that $u$ and $v$ are in the same column $C_{j}$, and respectively in rows $R_{i_{u}}$ and $R_{i_{v}}$. Suppose that neither $u^{\prime}$ nor $v^{\prime}$ is in $C_{j}$ (otherwise there are three of those vertices in $C_{j}$), and that one of them, without loss of generality $u^{\prime}$, is neither in $R_{i_{u}}$ nor in $R_{i_{v}}$. Thus $u^{\prime}$ is adjacent to both $u$ and $v$. Moreover, since $v^{\prime}$ cannot be both in $R_{i_{u}}$ and $R_{i_{v}}$, $v^{\prime}$ is adjacent to at least one of $u,v$, without loss of generality $v$. Then $uu^{\prime}vv^{\prime}$ is a path of $G_{n,m}$ with no center, a contradiction.
Therefore $u^{\prime},v^{\prime}$ belong to the union of the rows $R_{i_{u}}$ and $R_{i_{v}}$. If they are in the same row, then three of the vertices $u,v,u^{\prime},v^{\prime}$ are in a same row, so we may assume (w.l.o.g.) that $u^{\prime}$ belongs to $R_{i_{u}}$ and $v^{\prime}$ belongs to $R_{i_{v}}$. Suppose that $u^{\prime},v^{\prime}$ are not both in the same column: then $uv^{\prime}u^{\prime}v$ is a path of $G_{n,m}$ with no center, a contradiction.
@@
Suppose that there exist two distinct vertices $u,v\in V(G_{n,m})$ that are in a same column $C_{j}$ such that $\varphi(u)=\varphi(v)$. Let $R_{i_{u}}$ and $R_{i_{v}}$ be the two rows containing $u$ and $v$, respectively.
Observe that the graph $G^{\prime}=G_{n,m}-C_{j}$ contains exactly $nm-n$ vertices, and none of these vertices can be colored $\varphi(u)$, so if each vertex of $G^{\prime}$ has a unique color in $G^{\prime}$, then $\varphi$ uses at least $|V(G^{\prime})|+1>nm-n$ colors, a contradiction. Thus there exist two distinct vertices $x$ and $y$ in $G_{n,m}-C_{j}$ such that $\varphi(x)=\varphi(y)$. By Section 4.2, they are in rows $R_{i_{u}},R_{i_{v}}$, either both in one of them (then without loss of generality $R_{i_{u}}$) or one in each row, and both in the same column.
@@
Suppose that $G_{n,m}$ contains three distinct vertices $u$, $v$, and $w$ such that $\varphi(u)=\varphi(v)=\varphi(w)$. Since these three vertices are pairwise non-adjacent, they either all belong to the same row or to the same column. That contradicts Section 4.2. ∎
@@
Let $\{u_{k},v_{k}\}$, $k\in\{a,b,c\}$, be three disjoint pairs of vertices such that $\varphi(u_{k})=\varphi(v_{k})$ for all $k\in\{a,b,c\}$. Then, four vertices among those six belong to the same column or to the same row, and all the other $nm-6$ vertices have a unique color.
@@
Recall that $\varphi$ uses at most $nm-n$ colors. By Section 4.2, each of the colors appears at most twice. Let $k$ be the number of colors that appear twice. Then, the total number of colors used by $\varphi$ is $nm-k$, and, hence, $k\geqslant n$. On the other hand, $k\leqslant 3$, since $k\geqslant 4$ would contradict Section 4.2. Therefore, $n\leqslant k\leqslant 3$, and since $n\geqslant 3$, we infer that $m\leqslant n=k=3$. Since $k=3$, there exist three disjoint pairs of vertices, each of the same color. Hence, by Section 4.2, four vertices among those six belong to the same column or to the same row. However, this implies that $m\geqslant 4$ or $n\geqslant 4$, respectively, a contradiction. ∎
@@
It is easy to check that this is a linear coloring, however it is not a centered coloring, as there is no unique color. This graph is an induced subgraph of the complement $G_{2,4}$ of the $2\times 4$-rook’s graph. So any extension of the above coloring into a linear coloring of $G_{2,4}$ (for instance obtained by giving unique colors to the remaining vertices) is not a centered coloring.

### v2

@@ 3 General bounds
Following [KOP+21], Bose et al. [BDH+22] relaxed the notion of a $k\times k$ grid to a family of graphs called $k\times k$ pseudogrids, that have the following properties:
@@
    A graph $G$ contains $\boxplus_{k}$ as a minor if and only if $G$ contains a $k\times k$ pseudogrid as a subgraph (that is, $k\times k$ cor:treedeg are exactly the subgraph-minimal graphs containing $\boxplus_{k}$ as a minor).
@@
[[CNP21]] For every two integers $w,d>0$ and every graph $G$, if $\operatorname{\chi_{\mathrm{cen}}}(G)\geqslant\alpha wd$}, then either $\operatorname{\mathbf{tw}}(G)\geqslant w$ or $G$ contains as a subgraph a subcubic tree $T$ with $\operatorname{\chi_{\mathrm{cen}}}(T)\geqslant d$}.
@@
In the remaining outcome of Section 3, $G$ contains a subcubic tree $T$ with $\operatorname{\chi_{\mathrm{cen}}}(T)\geqslant d$ as a subgraph. We apply Section 2.2 and get $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant\operatorname{\chi_{\mathrm{lin}}}(T)\geqslant d/\log 3=\left\lfloor\sqrt{k/\alpha}\right\rfloor/\log 3$. So similarly as above, $k\leqslant c_{2}\operatorname{\chi_{\mathrm{lin}}}(G)^{2}$ for some constant $c_{2}>0$. Overall the claimed statement holds for $c=\max(c_{0},c_{1},c_{2})$. ∎
@@
$\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant c\cdot(\operatorname{\mathbf{tw}}(G)+1)\cdot\left(\operatorname{\chi_{\mathrm{lin}}}(G)+\frac{1}{\log 3}\right),$
@@
Treewidth does not increase when taking subgraphs, so only the second outcome of the theorem may hold. Hence, $G$ contains a subcubic tree $T$ with $\operatorname{\chi_{\mathrm{cen}}}(T)\geqslant d$.
@@
As a corollary (and using that $\operatorname{\chi_{\mathrm{cen}}}(G)=\operatorname{\chi_{\mathrm{lin}}}(G)$ when $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 2$, see Section 5.1) we get the following bound.
@@
$\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant c\cdot(\operatorname{\mathbf{tw}}(G)+1)\cdot\operatorname{\chi_{\mathrm{lin}}}(G)\,$,
@@
In graph classes where the treewidth is linearly bounded from above by the clique number (which is a lower bound for $\operatorname{\chi_{\mathrm{lin}}}$), Section 3 implies a quadratic bound. In particular, this is the case for intersection graphs of connected subgraphs of graphs of bounded treewidth. For a class $\mathcal{G}$ of graphs, let $I(\mathcal{G})$ be the class of region intersection graphs of graphs in $\mathcal{G}$, defined as follows. For $H\in\mathcal{G}$ and a family $\{H_{j}\}_{j\in J}$ of connected subgraphs of $H$, let $G$ be the graph with vertex set $J$ in which two distinct vertices $i$ and $j$ are adjacent if and only if $H_{i}$ and $H_{j}$ have a vertex in common. Region intersection graphs have been studied as a common generalization of many classes of geometric intersection graphs (see [LEE17]). For any graph class $\mathcal{G}$ of graphs with bounded treewidth, region intersection graphs of graphs in $\mathcal{G}$ have treewidth linearly bounded from above by the clique number. This follows from the following property, shown in the proof of [BGT98a, Lemma 2.4].
@@
[[BGT98a]] Let $k\geqslant 0$ be an integer and let $\mathcal{G}$ be a class of graphs with treewidth at most $k$. Then each graph $G\in I(\mathcal{G})$ has a tree decomposition in which each bag is a union of at most $k+1$ cliques.
@@
Let $k\geqslant 0$ be an integer and let $\mathcal{G}$ be a class of graphs with treewidth at most $k$. Then each graph $G\in I(\mathcal{G})$ satisfies $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant c(k+1)(\operatorname{\chi_{\mathrm{lin}}}(G)^{2}+\operatorname{\chi_{\mathrm{lin}}}(G)/\log 3)$, where $c=\alpha\log 3<3.7$.
@@
Let $G\in I(\mathcal{G})$. By Section 3, $G$ has a tree decomposition in which each bag is a union of at most $k+1$ cliques. Hence, $\operatorname{\mathbf{tw}}(G)\leqslant(k+1)\omega(G)-1$. By Section 3 and using the fact that $\omega(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)$ we infer the following.
$\operatorname{\chi_{\mathrm{cen}}}(G)$ $\leqslant c(\operatorname{\mathbf{tw}}(G)+1)(\operatorname{\chi_{\mathrm{lin}}}(G)+1/\log 3)$
$\leqslant c(k+1)(\operatorname{\chi_{\mathrm{lin}}}(G)^{2}+\operatorname{\chi_{\mathrm{lin}}}(G)/\log 3)\,.\qed$
{corollary}
For every class $\mathcal{G}$ of graphs with bounded treewidth, each graph $G\in I(\mathcal{G})$ satisfies $\operatorname{\chi_{\mathrm{cen}}}(G)=\operatorname{\mathcal{O}}(\operatorname{\chi_{\mathrm{lin}}}(G)^{2})$.
Since every chordal graph is an intersection graph of subtrees in a tree (see [BUN74, WAL78, GAV74]) and every circular-arc graph is an intersection graph of paths in a cycle, the cases $k\in\{1,2\}$ of Section 3 imply the following generalizations of the analogous result from [KOP+21] for interval graphs (Section 1).
@@
For every tree $T$, $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant c\cdot\operatorname{\chi_{\mathrm{lin}}}(T)$, where $c=\alpha\log 3<3.7\,.$
@@
4 When centered is linear or almost linear
Interesting cases of graphs satisfying Section 1 are the graphs for which the two chromatic numbers are equal. A sufficient condition for this is that every linear coloring of the graph is centered. Three graph classes with this property were already identified by Kun et al. [KOP+21]. In Section 4.1, we identify two more.
@@ 4 When centered is linear or almost linear
We conclude this preamble with definitions and a simple lemma that are used in several of the following subsections.
Given a positive integer $k$, a complete $k$-partite graph is any graph $G$ such that $V(G)$ admits a partition into $k$ nonempty parts such that two vertices of $G$ are adjacent if and only if they belong to different parts. A complete multipartite graph is any graph $G$ that is complete $k$-partite for some $k$. Note that if $G$ is complete multipartite, then $G$ is complete $k$-partite for a unique $k$; moreover, the corresponding partition into $k$ parts is unique. Complete multipartite graphs are cographs, hence, every linear coloring is centered (see [KOP+21]).
{lemma}
Proof.
@@ 4.1 Graphs for which every linear coloring is centered
Kun et al. [KOP+21] observed that this property holds for graphs with independence number at most two, cographs, and graphs with maximum degree at most $2$. The result for graphs with independence number at most two can be generalized as follows.
@@
$\overline{H}$ is connected. Since $\overline{H}$ is paw-free, by a result of Olariu [OLA88] $\overline{H}$ is either triangle-free or complete multipartite. In the first case the independence number of $H$ is at most $2$, so by [KOP+21, Lemma 1] $\varphi$ is a centered coloring, a contradiction. In the second case, as $H$ is connected, $H$ is a complete graph and the same conclusion follows.
@@
$\overline{H}$ is disconnected. Then $H$ has a bipartition $(A,B)$ of its vertex set into two nonempty sets with all edges between them. As $\varphi$ is a proper coloring, $\varphi$ colors $A$ and $B$ with disjoint sets of colors. In addition, by Section 4, in one of $A$ and $B$ $\varphi$ assigns different colors to all the vertices. We may assume without loss of generality that the vertices of $A$ get colored with different colors. Then, any vertex of $A$ is a center of $H$, contradicting the choice of $H$. ∎
A useful observation, also used in [KOP+21], is that the property holds for graphs such that every connected induced subgraph has a Hamiltonian path. {observation} If $G$ is such that every connected induced subgraph of $G$ has a Hamiltonian path, then every linear coloring of $G$ is centered. Indeed, if $H$ admits a linear coloring $\psi$, then for every connected subgraph $H$ of $G$, $V(H)$ induces a connected subgraph that has a Hamiltonian path $P_{H}$. Then $P_{H}$ admits a center, thus $H$ admits a center, and therefore $\psi$ is a centered coloring of $G$.
A graph $G$ is an interval graph if it is an intersection graph of a family of closed intervals on the real line. If, moreover, none of the intervals contains another, then $G$ is a proper interval graph. Kun et al. [KOP+21] showed that every interval graph $G$ satisfies $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)^{2}$. If $G$ is a proper interval graph, then $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$, as implied by the fact that every proper interval graph is (claw, net)-free [^3] The claw is the graph obtained by attaching 3 pendant vertices to a vertex, and the net is the graph obtained by attaching a pendant vertex to each vertex of a $K_{3}$. and the following. {proposition}If $G$ is (claw, net)-free, then every linear coloring of $G$ is centered. In particular, $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$.
@@ 4.2 Complete multipartite graphs and complements of rook’s graphs
{theorem}
Let $G$ be a complete multipartite graph. Then $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$, with both quantities equal to np+1n-p+1 where $n$ is the order of $G$ and $p$ is the maximum cardinality of a part of $G$.
@@
To see that $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant n-p+1$, observe that we can color the largest part $Y$ of $G$ with the same color and all the other vertices with different colors. This is a centered coloring since every subgraph of $G$ either contains a single vertex from $Y$, or contains a uniquely colored vertex from GYG-Y.
To show that $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant n-p+1$, consider a linear coloring $\varphi$ and let $k$ be the number of colors used. If all vertices of $G$ are colored with unique colors, then $k=n\geqslant n-p+1$. Hence, we may assume that two vertices $u,v$ receive the same color. Then $u$ and $v$ are non-adjacent and thus belong to the same part $X$ of $G$. Since the pair $(X,V(G)\setminus X)$ is a bipartition of $V(G)$ with all edges between $X$ and $V(G)\setminus X$, by Section 4, $\varphi$ assigns different colors to all the vertices in one of $X$ and $V(G)\setminus X$. Since $\varphi(u)=\varphi(v)$, we infer that $\varphi$ assigns different colors to all the vertices in $V(G)\setminus X$. Furthermore, since $\varphi$ is a proper coloring, no color of a vertex in $V(G)\setminus X$ can be used on any vertex in $X$. Hence, $k\geqslant n-|X|+1\geqslant n-p+1$, as claimed. We conclude that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant n-p+1\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)$, therefore, equalities must hold throughout. ∎
For all $n,m\geqslant 1$, the $n\times m$ rook’s graph is a graph on $nm$ vertices, arranged into $m$ columns with $n$ vertices each and into $n$ rows with $m$ vertices each, such that two vertices are adjacent if and only if they are in the same column or in the same row.
@@
Assume now that $m\geqslant 2$ and $n\geqslant 3$, with $n\geqslant m$. Since $G_{n,m}$ is a subgraph of the complete multipartite graph with $m$ parts of $n$ vertices, we obtain from Sections 2.2 and 4.2 that $\operatorname{\chi_{\mathrm{cen}}}(G_{n,m})\leqslant nm-n+1$; so let us show that $\operatorname{\chi_{\mathrm{lin}}}(G_{n,m})\geqslant nm-n+1$. Suppose for a contradiction that there exists a linear coloring $\varphi$ of $G_{n,m}$ with at most nmnnm-n colors.
@@
Suppose that $u$ and $v$ are in the same column $C_{j}$, and respectively in rows $R_{i_{u}}$ and $R_{i_{v}}$. Suppose that neither $u^{\prime}$ nor $v^{\prime}$ is in $C_{j}$ (otherwise there are three of those vertices in $C_{j}$), and that one of them, without loss of generality $u^{\prime}$, is neither in $R_{i_{u}}$ nor in $R_{i_{v}}$. Thus, $u^{\prime}$ is adjacent to both $u$ and $v$. Moreover, since $v^{\prime}$ cannot be both in $R_{i_{u}}$ and $R_{i_{v}}$, $v^{\prime}$ is adjacent to at least one of $u,v$, without loss of generality $v$. Then $uu^{\prime}vv^{\prime}$ is a path of $G_{n,m}$ with no center, a contradiction.
Therefore, $u^{\prime},v^{\prime}$ belong to the union of the rows $R_{i_{u}}$ and $R_{i_{v}}$. If they are in the same row, then three of the vertices $u,v,u^{\prime},v^{\prime}$ are in a same row, so we may assume (w.l.o.g.) that $u^{\prime}$ belongs to $R_{i_{u}}$ and $v^{\prime}$ belongs to $R_{i_{v}}$. Suppose that $u^{\prime},v^{\prime}$ are not both in the same column: then $uv^{\prime}u^{\prime}v$ is a path of $G_{n,m}$ with no center, a contradiction.
@@
Suppose that there exist two distinct vertices $u,v\in V(G_{n,m})$ that are in the same column $C_{j}$ such that $\varphi(u)=\varphi(v)$. Let $R_{i_{u}}$ and $R_{i_{v}}$ be the two rows containing $u$ and $v$, respectively.
Observe that the graph $G^{\prime}=G_{n,m}-C_{j}$ contains exactly nmnnm-n vertices, and none of these vertices can be colored $\varphi(u)$, so if each vertex of $G^{\prime}$ has a unique color in $G^{\prime}$, then $\varphi$ uses at least $|V(G^{\prime})|+1>nm-n$ colors, a contradiction. Thus there exist two distinct vertices $x$ and $y$ in $G_{n,m}-C_{j}$ such that $\varphi(x)=\varphi(y)$. By Section 4.2, they are in rows $R_{i_{u}},R_{i_{v}}$, either both in one of them (then without loss of generality $R_{i_{u}}$) or one in each row, and both in the same column.
@@
Suppose that $G_{n,m}$ contains three distinct vertices $u$, $v$, and $w$ such that $\varphi(u)=\varphi(v)=\varphi(w)$. Since these three vertices are pairwise non-adjacent, they either all belong to the same row or to the same column. This contradicts Section 4.2. ∎
@@
Let $\{u_{k},v_{k}\}$, $k\in\{a,b,c\}$, be three disjoint pairs of vertices such that $\varphi(u_{k})=\varphi(v_{k})$ for all $k\in\{a,b,c\}$. Then, four vertices among those six belong to the same column or to the same row, and all the other nm6nm-6 vertices have a unique color.
@@
Recall that $\varphi$ uses at most nmnnm-n colors. By Section 4.2, each of the colors appears at most twice. Let $k$ be the number of colors that appear twice. Then, the total number of colors used by $\varphi$ is nmknm-k, and, hence, $k\geqslant n$. On the other hand, $k\leqslant 3$, since $k\geqslant 4$ would contradict Section 4.2. Therefore, $n\leqslant k\leqslant 3$, and since $n\geqslant 3$, we infer that $m\leqslant n=k=3$. Since $k=3$, there exist three disjoint pairs of vertices, each of the same color. Hence, by Section 4.2, four vertices among those six belong to the same column or to the same row. However, this implies that $m\geqslant 4$ or $n\geqslant 4$, respectively, a contradiction. ∎
@@
It is easy to check that this is a linear coloring; however, it is not a centered coloring, as there is no unique color. This graph is an induced subgraph of the complement $G_{2,4}$ of the $2\times 4$-rook’s graph. So any extension of the above coloring into a linear coloring of $G_{2,4}$ (for instance, obtained by giving unique colors to the remaining vertices) is not a centered coloring.